OpenStudy (anonymous):

rewrite the ff.in the of Y=mx+b 2x-4y+10=0 2x-3y+6=0

OpenStudy (crashonce):

well move y to one side and everything to the other so: y=-(-2x-10)/4 = x/2+5/2 y=-(-2x-6) /3 = 2x/3 + 2

OpenStudy (crashonce):

@scale22 do you understand?

OpenStudy (anonymous):

uh uh. i don't.

OpenStudy (crashonce):

first you move everythin except the y to one side

OpenStudy (crashonce):

then you simplify the RHS

OpenStudy (anonymous):

can you give the answer. all i need is the answer.?

OpenStudy (crashonce):

y= x/2+5/2 y= 2x/3 + 2 but i need you to learn @scale22 i hope you understand anyway

OpenStudy (anonymous):

i have my answer. i just wanna check it. if its correct.